# Rule-based Generation of de Bruijn Sequences: Memory and Learning

### v1

@@ 4 Characterization of de Bruijn rules
As stated in the previous section, certain generating rules produce sequences with maximum period—we refer to them as de Bruijn sequences. Therefore, a complete characterization of such rules emerges as a fundamental goal. The following properties help identify and constrain the set of de Bruijn rules, and can be used to systematically search for them:
@@
  • Boundary Conditions: The binary representations of de Bruijn rules must start with 00 and end with $1$. This condition arises from the convention in rule ordering, which arranges input strings from the highest binary value ($11\ldots 1$ 11) to the lowest ($00\ldots 0$ 00), and from the necessity to avoid fixed points (e.g., if $R(11\ldots 11)=1$, the sequence remains constant). This constraint immediately reduces the number of candidate rules by a factor of 4. For example, for $\mu=4$, the number of potentially valid rules is reduced from $655362^{2^{4}}=$ 65536 to $16384$.
@@
  • Symmetry and Parity: The binary representations of de Bruijn rules are symmetric with respect to their midpoint, such that each half is the complement of the other. This symmetry ensures parity: the number of 00s equals the number of $1$s, although the balance may be broken within each half. Under this constraint, valid de Bruijn rules correspond to multiples of certain numbers derived from sequences of Evil Odd numbers (A129771, ), multiplied by factors denoted as $\phi(\mu)$.
    For instance, if we denote the Evil Odd Numbers, i.e., an odd number whose binary representation contains an even number of $1$, for each $\mu$ as $R(\mu)_{k}$, being $k=1,2,\ldots,C(\mu)$, we get:
    $R(2)_{1}=1\times 3=3,\quad R(3)_{1}=3\times 15=45,\quad R(3)_{2}=5\times 15=75$
    The factors follow the recursive relation:
    $\phi(\mu+1)=\phi(\mu)\cdot\left(2^{2^{\mu}}+1\right)$ (5)
@@
    For instance, $\phi(2)=1$, $\phi(3)=15$, $\phi(4)=255$, and $\phi(5)=65535$. Table 2 presents all de Bruijn rules for $\mu=4$ whose decimal representations are products of an Evil Odd number and the corresponding $\phi(\mu)$. Note, however, that not all Evil Odd numbers yield valid de Bruijn rules, and the problem of determining which ones do remains open for larger values of $\mu$. For $\mu=4$, applying these conditions reduce the feasible set of de Bruijn rules to 32.
@@
  • Constrained Position Pairs: There exist pairs of positions in the binary representation of de Bruijn rules that cannot simultaneously take the same value (1 for even $\mu$-values and 0 for odd ones). This condition is applied only to the first half of the binary string for symmetry (according to the previous item) and depends on $\mu$ according to a recursive pattern. Specifically: For $\mu=2$, the constrained positions are $p^{1}_{2}=1$ and $p^{2}_{2}=2\,p^{1}_{2}$, and
@@
    • If $\mu=4k$, then $p^{1}_{4k}=p^{2}_{4k-1}$ and $p^{2}_{4k}=2\,p^{1}_{4k}$,
    for $k=1,2,\ldots$. This structural condition eliminates $1/4$ of the remaining candidates. For example, for $\mu=4$ these positions are $p^{1}_{4}=3$ and $p^{2}_{4}=6$. For $\mu=5$, $p^{1}_{5}=6$ and $p^{2}_{5}=11$.
    Remarkably, for $\mu=4$, the final count of feasible de Bruijn rules after applying all constraints is 24. In general, for all values of $\mu$, the feasible set exceeds the actual number of de Bruijn rules (see Table 3).
  • Symmetric Rule Invariance: If a rule of the form:
@@
    is a de Bruijn rule, then its mirrored version:
[…]
@@
The application of the properties stated in the previous paragraphs to the entire set of generating rules significantly reduces the number of feasible rules that can yield de Bruijn sequences. Table 3 presents the total number of rules, the number of de Bruijn sequences, the number of feasible rules after applying the constraints, and the corresponding ratios.
As can be observed, the number of feasible rules that need to be checked to find de Bruijn rules is drastically smaller than the total number of rules for each $\mu$ value. For instance, when $\mu=6$, the reduction factor is approximately $10^{-11}$. In the same case, the ratio between the number of rules that actually yield de Bruijn sequences and the number of feasible rules is around $0.17$. For such low values of $\mu$, a brute-force approach may still be used to generate all de Bruijn sequences. However, a more efficient and structured methodology will be presented in the next section.
@@
$\mu$ $C(\mu)$#\## Feasible ##\## de Bruijn Feasible/Total de Bruijn/Total de Bruijn/Feasible
@@
4 65536 24 16 0.000366211 0.000244141 0.66666667
5 4294967296 6144 2048 1.4305E-06 4.7683E-07 0.33333333
6 1.84467E+19 402653184 67108864 2.1827E-11 3.6379E-12 0.16666667
7 3.40282E+38 1.7293E+18 1.4411E+17 5.082E-21 4.2351E-22 0.08333333
8 1.15792E+77 3.1901E+37 1.3292E+36 2.7550E-40 1.1479E-41 0.04166667
9 1.3408E+154 1.0855E+76 2.2615E+74 8.0964E-79 1.6867E-80 0.02083333
@@ 5 Neural Networks to classify de Bruijn rules
An alternative approach to identifying de Bruijn rules is to apply machine learning methods. In particular, neural networks are especially well-suited for classification tasks Bishop . In this section, we present a neural network model for classifying the feasible rules into two categories: de Bruijn rules (coded by 1) and the rest (0). Although classification based on the sequence period is also possible, it would require accounting for dependence on initial conditions. The classification is performed for $\mu=5$ and $\mu=6$ (see Table 4).
We implemented a binary classification model using a feedforward neural network in R Rcore . The analysis followed these main steps:
@@
    Dataset splitting: The data was randomly split into training (80%) and testing (20%) subsets to evaluate model performance on unseen data.
@@
    Model specification and training: A feedforward neural network was constructed using the keras package with a TensorFlow backend Allaire ; Rcore .
@@
For the case $\mu=5$, the neural network architecture includes:
  • An input layer with 14 features (bits),
  • A hidden dense layer with 32 units and ReLU activation,
  • A second hidden dense layer with 16 units and ReLU activation,
@@
The model was compiled using the Adam optimizer (learning rate $=0.001$), binary cross-entropy loss function, and accuracy as a performance metric. The dataset includes the complete set of 6144 feasible rules, one-third of which are de Bruijn rules (see Table 3). Training was performed over 100 epochs with a batch size of 4. Model evaluation was conducted on the test set by calculating accuracy, sensitivity, and specificity (see Table 4). Class labels were assigned using a threshold of 0.5 on the predicted probabilities. The resulting classifier achieved outstanding performance, with an accuracy exceeding 99% for $\mu=5$.
The more challenging case of $\mu=6$ involves a rule space of size on the order of $10^{19}$, with approximately 67 million de Bruijn rules (see Table 3). As a matter of fact, we can only use a sample of the total rule space that is randomly analyzed and classified into the two classes.
@@
  • Three hidden dense layers with 64, 64, and 8 units respectively,
  • ReLU activations in all hidden layers,
@@
Training used a batch size of 64 and a learning rate of $0.001$. The dataset consists of a sample of $2\times 10^{6}$ feasible rules, half of which are de Bruijn rules. As before, 80% of the dataset was used for training, and the remainder for testing. A threshold of 0.5 was again applied to the output probabilities to assign class labels. As shown in Table 4, this model also demonstrated excellent classification metrics.
Table 4: Evaluation metrics for the neural network models applied to classify the de Bruijn rules for memories with $\mu=5$ and $\mu=6$. Total refers to the sum: $TP+FP+TN+FN$, and corresponds to one fith of the whole data set. Positive class considered: 1 $\to$ de Bruijn.
Metric (definition) $\mu=5$ $\mu=6$
True Positives (TP) 397 198563
False Positives (FP) 3 19839
True Negatives (TN) 820 180668
@@
Balanced Accuracy, (Sens. + Spec.) / 2 0.9872 0.9482
Detection Rate, TP / Total 0.3230 0.4976
Detection Prevalence, (TP + FP) / Total 0.3255 0.5473
True Prevalence, (TP + FN) / Total 0.3304 0.4999
@@ 6 Discussion
The successive application of updating rules to an initial configuration of $\{0,1\}$ generates a binary sequence that becomes asymptotically periodic. The size of the initial configuration depends on the memory of the rule, denoted by $\mu$, which is the number of digits required to update the next one. For large values of $\mu$, the number of possible generating rules becomes so large that an exhaustive analysis of the resulting patterns is infeasible. Particularly relevant is a specific and highly constrained subset of rules, that we named as de Bruijn rules, that generate sequences of maximum period, known as de Bruijn sequences. Although these sequences represent only a tiny fraction of the entire set of possible sequences, their sheer number for large $\mu$ remains enormous, and the complete and effective generation of all such sequences is still an open problem.

### v2

@@ 4 Characterization of de Bruijn Rules
As stated in the previous section, certain generating rules produce sequences with maximum period—we refer to them as de Bruijn rules. Therefore, a complete characterization of such rules emerges as a fundamental goal. The following properties help identify and constrain the set of de Bruijn rules and can be used to systematically search for them:
@@
  • Boundary Conditions: The binary representations of de Bruijn rules must start with 0 and end with $1$. This condition arises from the convention in rule ordering, which arranges input strings from the highest binary value ($11\ldots 11$) to the lowest ($00\ldots 00$) and from the necessity to avoid fixed points (e.g., if $R(11\ldots 11)=1$, the sequence remains constant). This constraint immediately reduces the number of candidate rules by a factor of 4. For example, for $\mu=4$, the number of potentially valid rules is reduced from $2^{2^{4}}$ = 65,536 to 16,384.
@@
  • Symmetry and Parity: The binary representations of de Bruijn rules are symmetric with respect to their midpoint, such that each half is the complement of the other. This symmetry ensures parity: the number of 0s equals the number of $1$s, although the balance may be broken within each half. Under this constraint, valid de Bruijn rules correspond to multiples of certain Numbers derived from sequences of Evil Odd numbers (A129771, ), multiplied by factors denoted as $\phi(\mu)$ (see Appendix A). Let $R_{\mu}$ be a rule for a given $\mu$. Then,
    $R_{\mu}=\phi(\mu)\,(A+1)$ (5)
    In Appendix A, we show that the factor $\phi(\mu)$ is
    $\phi(\mu)=2^{2^{\mu-1}}-1$ (6)
    For instance, $\phi(2)=3$, $\phi(3)=15$, $\phi(4)=255$, and $\phi(5)$ = 65,535. The consideration of this property causes a huge reduction in the feasible set of the de Bruijn rules, from 16,384 to 64 for the case $\mu=4$.
  • Evil Odd Number divisibility: It is also empirical evidence that factor $A+1$ in Equation (5) must be an Evil Odd Number for $R_{\mu}$ to be a de Bruijn rule. For each $\mu$, there is a unique $\phi$ that divides $R_{\mu}$, which results in multiple remainders that are Evil Odd Numbers. To consider that, it is convenient to denote this factor as $A(\mu)_{k}$, representing the Evil Odd Numbers for $k=1,2,\ldots,C(\mu)$. For instance,
    $A(2,1)=1\times 3=3,\quad A(3,1)=3\times 15=45,\quad A(3,2)=5\times 15=75$
@@
    Table 2 presents all de Bruijn rules for $\mu=4$ whose decimal representations are products of an Evil Odd Number and the corresponding $\phi(\mu)$. Note, however, that not all Evil Odd Numbers yield valid de Bruijn rules, and the problem of determining which ones do remains open for larger values of $\mu$. For $\mu=4$, applying these conditions reduce the feasible set of de Bruijn rules to 32.
@@
  • Constrained Position Pairs: There exist pairs of positions in the binary representation of de Bruijn rules that cannot simultaneously take the same value (1 for even $\mu$-values and 0 for odd ones). This condition is applied only to the first half of the binary string for symmetry (according to the previous item) and depends on $\mu$ according to a recursive pattern.
    Let $p_{\mu}^{j}$ be the binary value of the $j$ position in a de Bruijn rule of memory $\mu$. For $\mu=2$, the constrained positions are $p^{1}_{2}=1$ and $p^{2}_{2}=2\,p^{1}_{2}$, and
@@
    • If $\mu=2\,(k+1)$, then $p^{1}_{2\,(k+1)}=p^{2}_{2\,k+1}$ and $p^{2}_{2\,(k+1)}=2\,p^{1}_{2\,(k+1)}$.
    for $k=1,2,\ldots$.
    This structural condition eliminates $1/4$ of the remaining candidates. For example, for $\mu=3$, these positions are $p^{1}_{3}=2$ and $p^{2}_{3}=3$. For $\mu=4$, $p^{1}_{4}=3$ and $p^{2}_{4}=6$, and for $\mu=5$, $p^{1}_{5}=6$ and $p^{2}_{5}=11$.
    Remarkably, for $\mu=4$, the final count of feasible de Bruijn rules after applying all constraints is 24. In general, for all values of $\mu>3$, the feasible set exceeds the actual number of de Bruijn rules (see Table 3).
  • Symmetric Rule Invariance: If a rule of the form
@@
    is a de Bruijn rule, then its mirrored version
    $0\,a_{k}\,a_{k-1}\,\ldots\,a_{2}\,a_{1}\,0\,1\,1-a_{k}\,1-a_{k-1}\,\ldots\,1-a_{2}\,1-a_{1}\,1$
@@
Let us summarize the reduction procedure that results from the application of these properties for the case $\mu=4$. There are 65,536 rules that are filtered as follows:
  • After applying the boundary conditions, 16,384 rules remain feasible.
  • Applying symmetry, i.e., dividing by $\phi(4)=255$, 64 rules remain.
  • The requirement that the remainder is an Evil Odd Number reduces the previous number to 32 rules.
  • Verifying the constrained position pairs reduces the feasible set to 24 rules.
From these 24 feasible rules, only 16 are de Bruijn (see Table 2).
The application of the properties stated in the previous paragraphs to the entire set of generating rules significantly reduces the number of feasible rules that can yield de Bruijn sequences (an example can be found in Appendix C). Table 3 presents the total number of rules, the number of de Bruijn sequences, the number of feasible rules after applying the constraints, and the corresponding ratios. As can be observed, the number of feasible rules that need to be checked to find de Bruijn rules is drastically smaller than the total number of rules for each $\mu$ value. For instance, when $\mu=6$, the reduction factor is approximately $10^{-11}$. In the same case, the ratio between the number of rules that actually yield de Bruijn sequences and the number of feasible rules is around $0.17$. For such low values of $\mu$, a brute-force approach may still be used to generate all de Bruijn sequences. However, a more efficient and structured methodology will be presented in the next section.
@@
$\mu$ $C(\mu)$ #\# Feasible #\# de Bruijn Feasible/Total de Bruijn/Total de Bruijn/Feasible
@@
4 65,536 24 16 0.000366211 0.00024414 0.66666667
5 4,294,967,296 6144 2048 1.4305 $\times$ 106-6 4.7683 $\times$ 107-7 0.33333333
6 1.84467 $\times$ 101919 402,653,184 67,108,864 2.1827 $\times$ 1011-11 3.6379 $\times$ 1012-12 0.16666667
7 3.40282 $\times$ 103838 1.7293 $\times$ 101818 1.4411 $\times$ 101717 5.0820 $\times$ 1021-21 4.2351 $\times$ 1022-22 0.08333333
8 1.15792 $\times$ 107777 3.1901 $\times$ 103737 1.3292 $\times$ 103636 2.7550 $\times$ 1040-40 1.1479 $\times$ 1041-41 0.04166667
9 1.3408 $\times$ 10154154 1.0855 $\times$ 107676 2.2615 $\times$ 107474 8.0964 $\times$ 1079-79 1.6867 $\times$ 1080-80 0.02083333
@@ 5 Neural Networks to Classify de Bruijn Rules
An alternative approach to identifying de Bruijn rules is to apply machine learning methods. In particular, neural networks are especially well suited for classification tasks Bishop . In this section, we present a neural network model for classifying the feasible rules into two categories: de Bruijn rules (coded by 1) and the rest (0). Although classification based on the sequence period is also possible, it would require accounting for dependence on initial conditions. The classification is performed for $\mu=5$ and $\mu=6$ (see Table 4).
We implemented a binary classification model using a feedforward neural network in R (version 4.5.0) Rcore . The analysis followed these main steps:
@@
    Dataset splitting: The data were randomly split into training (80%) and testing (20%) subsets to evaluate model performance on unseen data. The validation set was chosen to be $20\%$ of the training set in all cases.
@@
    Model specification and training: A feedforward neural network was constructed using the keras package (version 2.15.0) with a TensorFlow (version 2.16.0) backend Allaire ; Rcore .
@@
For the case $\mu=5$, the neural network architecture includes the following:
  • An input layer with 14 features (bits);
  • A hidden dense layer with 32 units and ReLU activation;
  • A second hidden dense layer with 16 units and ReLU activation;
@@
The model was compiled using the Adam optimizer (learning rate =0.001=0.001), binary cross-entropy loss function, and accuracy as a performance metric. The dataset includes the complete set of 6144 feasible rules, one-third of which are de Bruijn rules (see Table 3). Training was performed over 100 epochs with a batch size of 4. Model evaluation was conducted on the test set by calculating accuracy, sensitivity, and specificity (see Table 4). Class labels were assigned using a threshold of 0.5 on the predicted probabilities. The resulting classifier achieved outstanding performance, with an accuracy exceeding 99% for $\mu=5$.
The more challenging case of $\mu=6$ involves a rule space of size on the order of $10^{19}$, with approximately 67 million de Bruijn rules (see Table 3). As a matter of fact, we can only use a sample of the total rule space that is randomly analyzed and classified into the two classes. To get this sample of the feasible rule space, we used brute force to generate 1 million de Bruijn rules. In parallel, we generated a larger number of non-de Bruijn feasible rules and randomly selected 1 million of them. Note that the ratio of de Bruijn to feasible rules is about 0.17 (Table 3), which means that approximately 6 million feasible rules must be generated by brute force to obtain 1 million de Bruijn rules.
@@
  • Three hidden dense layers with 64, 64, and 8 units, respectively;
  • ReLU activations in all hidden layers;
@@
Training used a batch size of 64 and a learning rate of $0.001$. The dataset consists of a sample of $2\times 10^{6}$ feasible rules, half of which are de Bruijn rules. As before, 80% of the dataset was used for training, with 20% used for validation and the remainder for testing. A threshold of 0.5 was again applied to the output probabilities to assign class labels. As shown in Table 4, this model also demonstrated excellent classification metrics.
Table 4: Evaluation metrics for the neural network models applied to classify the de Bruijn rules for memories with $\mu=5$ and $\mu=6$. Total refers to the sum TP+FP+TN+FN and corresponds to one fifth of the whole dataset. The positive class considered 1 $\to$ de Bruijn.
Metric, Definition $\mu=5$ $\mu=6$
True Positives (TP) 397 198,563
False Positives (FP) 3 19,839
True Negatives (TN) 820 180,668
@@
Balanced Accuracy, (Sens. + Spec.)/2 0.9871 0.9482
Detection Rate, TP/Total 0.3230 0.4964
Detection Prevalence, (TP + FP) / Total 0.3255 0.5460
True Prevalence, (TP + FN) / Total 0.3303 0.4987
@@ 6 Discussion
The successive application of updating rules to an initial configuration of $\{0,1\}$ generates a binary sequence that becomes asymptotically periodic. The size of the initial configuration depends on the memory of the rule, denoted by $\mu$, which is the number of digits required to update the next one. For large values of $\mu$, the number of possible generating rules becomes so large that an exhaustive analysis of the resulting patterns is infeasible. Particularly relevant is a specific and highly constrained subset of rules that we named as de Bruijn rules and that generate sequences of maximum period, known as de Bruijn sequences. Although these sequences represent only a tiny fraction of the entire set of possible sequences, their sheer number for large $\mu$ remains enormous, and the complete and effective generation of all such sequences is still an open problem.